OpenStudy (anonymous):

what are some of codes that you know from the new extintion!!! dying to know i want to try them all!!!!

eclipsedstar (eclipsedstar):

What's a "new extinction"? xD

OpenStudy (anonymous):

extension is what i ment

imqwerty (imqwerty):

download nd run :P

eclipsedstar (eclipsedstar):

You mean @Jaynator495 's extension?

imqwerty (imqwerty):

no i mean the thing which i posted rn there ^

eclipsedstar (eclipsedstar):

Nope, I was talking to animalfan lol

jaynator495 (jaynator495):

animal this is considered spam, well... it will be until its in openstudy full time... xD

imqwerty (imqwerty):

:)

eclipsedstar (eclipsedstar):

It's still in a os-ish sense...so not exactly spam.

imqwerty (imqwerty):

try downloadin the thing i posted :P

OpenStudy (anonymous):

is the thing you posted a computer virus @imqwerty?

OpenStudy (anonymous):

because my computer antivirus/anti malware program automatically blocked the download

OpenStudy (anonymous):

@imqwerty

OpenStudy (anonymous):

i did and it said it will harm my computer so it canceld out by force...

imqwerty (imqwerty):

@heretohelpalways that is not a virus and @animalfan1 it will not harm ur pc...that dialogue box appears only because the file is of the format .vbs

jaynator495 (jaynator495):

Meh, sorta kinda a virus... he actually used a site to generate a window with text when you ran the script... X=MsgBox("Error while opening computer.Do you want to fix it?",4+64,"Computer") X=MsgBox("Unable to fix this error. Do you want to scan your Computer",3+48,"Title") X=MsgBox("Alert Virus has been detected. Do you want to delete this virus?",3+16,"Title") X=MsgBox("Unable to delete this virus ",1+64,"Critical Error") X=MsgBox("Virus is Activated",2+16,"Virus Alert") X=MsgBox("Deleting System Files",2+16,"File Deletion") X=MsgBox("Virus is copying your password",2+48,"Virus Alert") X=MsgBox("Please wait. Uploading your files to hack server. Do you want to stop it?",4+64,"File Transfer") X=MsgBox("Cannot stop. File transfer completed",1+16,"Completed") X=MsgBox("Your computer is hacked",1+64,"Hacked") X=MsgBox("You cannot Access your Computer Anymore",1+48,"Computer Hacked") X=MsgBox("This was a PRANK, there is no virus - imqwerty xD",1+64,"No Virus")

jaynator495 (jaynator495):

Dude, thats kinda pathetic... if your gonna do it do it right imqwerty xD

jaynator495 (jaynator495):

X=MsgBox("Alert Virus has been detected. Do you want to delete this virus?",3+16,"Title") What to lazy to make a title? ._.

imqwerty (imqwerty):

L jay... I mi forgot to put that title nd i don use any sites to do this :P

jaynator495 (jaynator495):

cummon, dun lie, we both know you dont know how to code x'D